### v1

# PTB-Image: A Scanned Paper ECG Dataset for Digitization and Image-based Diagnosis

Abstract
Electrocardiograms (ECGs) recorded on paper remain prevalent in clinical practice, yet their use presents challenges for automated analysis and digital storage. To address this issue, we introduce PTB-Image, a dataset comprising scanned paper ECGs with corresponding digital signals, enabling research on ECG digitization. We also provide VinDigitizer, a digitization baseline to convert paper-based ECGs into digital time-series signals. The method involves detecting signal rows, extracting waveforms from the background, and reconstructing numerical values from the digitized traces. We applied VinDigitizer to 549 scanned ECGs and evaluated its performance against the original PTB dataset (modified to match the printed signals). The results achieved a mean signal-to-noise ratio (SNR) of 0.01 dB, highlighting both the feasibility and challenges of ECG digitization, particularly in mitigating distortions from printing and scanning processes. By providing PTB-Image and baseline digitization methods, this work aims to facilitate advancements in ECG digitization, enhancing access to historical ECG data and supporting applications in telemedicine and automated cardiac diagnostics.
**footnotetext: These authors contributed equally to this work.
Keywords electrocardiogram  $\cdot$ digitization  $\cdot$ ecg diagnosis
1 Introduction
Electrocardiograms (ECG) serve as a fundamental tool in the detection and diagnosis of cardiovascular disease (CVD), which continues to be one of the leading causes of mortality worldwide. Early and accurate identification of CVD is critical for timely medical intervention, and recent advancements in algorithmic interpretation, particularly those leveraging deep learning techniques [1, 2, 3, 4, 5, 6, 7], have significantly improved the accessibility and accuracy of ECG-based diagnostics. However, these AI-driven approaches primarily rely on digital ECG data, which limits their applicability in real-world clinical settings where paper-based ECGs remain prevalent.
Despite the increasing availability of digital ECG systems, paper ECGs continue to be widely used due to their historical adoption, cost-effectiveness, and compatibility with legacy healthcare infrastructures. The continued reliance on paper ECGs presents significant challenges, as it restricts automated data analysis, complicates electronic health record (EHR) integration, and limits the development of robust AI models trained on diverse ECG data. Consequently, the digitization of paper ECGs is essential for enhancing data accessibility, facilitating large-scale retrospective studies, and reducing the costs associated with CVD diagnosis. Furthermore, an effective digitization pipeline would play a crucial role in expanding telemedicine capabilities, enabling remote diagnosis, and supporting AI-driven advancements in ECG interpretation.
However, the digitization of paper-based ECGs poses several technical challenges, including preserving signal fidelity, handling variations in paper quality and scanning artifacts, and ensuring that the digitization process remains scalable and generalizable across different healthcare settings. Addressing these challenges is necessary to prevent information loss during conversion and to maintain the accuracy of downstream diagnostic applications.
Recent efforts have been made to develop methodologies for digitizing paper-based ECG records [8, 9, 10, 11, 12, 13, 14, 15], with particular momentum driven by the PhysioNet Challenge 2024 [16]. This challenge utilizes a specialized software tool to generate synthetic images of scanned paper ECGs [17], offering a substantial dataset for training and evaluation purposes. However, while these synthetic ECG images provide a useful benchmark, they do not fully replicate the complexity and variability of real-world paper ECG scans, leading to potential overfitting when models trained on synthetic data are applied to actual clinical ECG records.
To bridge this gap, we introduce PTB-Image, a comprehensive dataset of scanned paper ECGs paired with their corresponding digital signals. This dataset is based on the PTB dataset [18], and designed to facilitate the development and evaluation of robust ECG digitization methods.
Key contributions:
  • We present PTB-Image, a realistic dataset of paper-scanned ECGs with corresponding digital signals, enabling the advancement of digitization techniques.
  • We provide baseline methods for ECG digitization and automated diagnosis using the dataset, serving as reference methodologies for future research and improvement.
2 The PTB-Image Dataset
The original PTB dataset [18] was collected using a non-commercial PTB prototype recorder with 16 input channels (14 ECG, 1 respiration, 1 line voltage), an input voltage range of ±16 mV (offset compensation up to ±300 mV), 100 Ohm input resistance, and 16-bit resolution (0.5 uV/LSB). Signals were synchronously sampled at 1000 Hz across a 0–1 kHz bandwidth, with noise levels reaching a maximum of 10 uV (pp) or 3 uV (RMS). The dataset comprises 549 records from 290 subjects (ages 17–87, mean 57.2), each represented by one to five recordings. Each record includes 15 simultaneous ECG signals: the standard 12-lead ECG (I, II, III, aVR, aVL, aVF, V1–V6) and the three Frank leads (Vx, Vy, Vz), digitized with 16-bit resolution over a ±16.384 mV range. Higher sampling rates (up to 10 kHz) may be available upon request. Most records include a .hea file containing clinical summaries with demographic information, diagnoses, and, where applicable, medical history, prescribed medications, coronary artery pathology, ventriculography, echocardiographic findings, and hemodynamic data.
In this work, we printed 12 out of the 15 leads (I, II, III, aVR, aVL, aVF, V1–V6) from each record in the original PTB dataset onto ECG paper using the ECG Image Kit [17]. Given the length of the original ECG signals, only a portion of each signal was printed. Consequently, the original digital records (.dat format) were modified to align with the printed segments, ensuring a valid basis for comparison. After printing, all 549 paper ECGs were scanned to complete the dataset collection. Figure 1 illustrates an example of a scanned ECG, with patient data anonymized.
Figure 1: Sample scanned paper ECG
The dataset is available HERE.
3 PTB-Image for ECG Digitization
3.1 Task Description
The digitization task aims to convert paper-scanned ECGs into digital time-series signals, enabling automated analysis and integration with electronic health records. Standard paper ECGs typically display all 12 leads in 2.5-second segments, arranged in three rows of four leads each. Additionally, an extended 10-second strip of lead II, V1, V2, or V5 is often included at the bottom (see Figure 1). The objective is to accurately extract and reconstruct the 12-lead ECG signals from the scanned images, preserving signal morphology and ensuring compatibility with digital processing techniques. The final output consists of digitally reconstructed 12-lead ECG signals, which can be further utilized for clinical evaluation and algorithmic interpretation.
3.2 Methodology
In this work, we use VinDigitizer [10], our approach in the Physionet Challenge 2024, for baseline evaluation. VinDigitizer follows a three-stage pipeline, which consists of detecting signal rows on the paper ECG, extracting the signal from the background, and assigning numerical values to each signal point.
The first stage focuses on isolating the sections of the image that contain ECG lead signals. Standard paper ECGs typically display all 12 leads in 2.5-second segments, arranged in three rows of four leads each. To facilitate this segmentation, a YOLOv8 [19] model was trained on a dataset of 300 annotated paper ECGs, enabling it to accurately detect and localize the four signal-containing rows.
Figure 2: Signal extraction procedure.
The second stage involves extracting numerical representations of the waveform within each segmented ECG row. The signal extraction process, illustrated in Figure 2, begins with converting the RGB image to grayscale, allowing for pixel intensity quantification. The grayscale image is then binarized using Otsu’s thresholding [20], where pixels with intensity values below the computed threshold are assigned a value of 0 (black), while those above the threshold are set to 255 (white).
Following binarization, the background grid is iteratively removed using the method outlined in [21]. The vertical and horizontal grid lines are detected based on column-wise white-point density. If the grid remains visible in the binary image, the grayscale-to-binary conversion threshold is reduced by 5%, and the process is repeated until the grid is no longer detectable, resulting in a clean, signal-only binary image.
In the final stage, numerical signal values are extracted from the processed binary image. This step follows the approach proposed in [21], where vertical white pixel clusters are treated as candidate signal nodes, and the waveform is reconstructed by identifying the least-cost path from left to right across the image. The algorithm first detects candidate nodes in each column and then evaluates possible connections to preceding non-empty columns. The cost of each connection is computed based on Euclidean distance and angular alignment, ensuring smooth waveform reconstruction. The path with the lowest total cost is selected as the extracted signal, and an array of vertical pixel coordinates along this path is returned as the final digital representation.
3.3 Evaluation Metrics
For the digitization task, we utilized the following metrics: Signal-to-Noise Ratio (SNR), SNR median, Kolmogorov-Smirnov (KS) metric, weighted absolute difference (WAD), and adaptive signed correlation index (ASCI).
Specifically, SNR (in dB) is defined as the ratio of the power of the original signal to the power of the noise, which is the difference between the original and the digitized signal:
$\text{SNR}=10\log_{10}\left(\frac{P_{\text{original}}}{P_{\text{original}-\text{digitized}}}\right)$ (1)
A higher SNR indicates that the digitized signal is closer to the original. If the digitized signal is identical to the original, the SNR is infinity.
3.4 Results
VinDigitizer was applied to all 549 scanned ECG files from the PTB-Image dataset to convert the printed waveforms back into digital signals. To assess the accuracy of the digitization process, we compared the extracted signals with the corresponding original PTB samples, which had been modified to match the specific portions of the signals that were printed on paper. This ensured a fair and precise evaluation, as the reference signals were aligned with the available printed segments. The signal reconstruction quality was quantified using the signal-to-noise ratio (SNR), yielding a mean SNR of 0.01 dB across the dataset. This result highlights the challenges associated with extracting clean signals from scanned paper ECGs, including variations in printing quality, scanner-induced distortions, and grid removal artifacts. Despite these challenges, the digitized signals maintained recognizable morphological features, demonstrating the feasibility of the proposed method for recovering ECG waveforms from paper records. Figure 3 presents an example of a digitized ECG signal alongside its original reference.
Figure 3: Digitization result of Lead II, patient 001, ID s0010_re.
4 Conclusion
In this work, we introduced PTB-Image, a dataset of scanned paper ECGs paired with their corresponding digital signals, to facilitate research on ECG digitization. To address the challenges of converting printed ECG waveforms back into digital form, we proposed VinDigitizer, a three-stage pipeline comprising signal row detection, background removal, and waveform extraction. Our method was applied to 549 scanned ECGs from PTB-Image, and its performance was evaluated by comparing the reconstructed signals with their original digital counterparts. The results, with a mean SNR of 0.01 dB, highlight both the feasibility and challenges of digitizing paper ECGs, particularly in terms of maintaining signal fidelity amidst printing and scanning artifacts.
Despite its effectiveness, the current approach has limitations, such as potential distortions introduced during the printing and scanning process, which may affect signal reconstruction accuracy. Future work could explore advanced deep-learning models for robust waveform extraction, improved preprocessing techniques for noise reduction, and domain adaptation methods to enhance generalizability across different paper ECG formats. Additionally, expanding the dataset to include ECGs from various sources and printing conditions would further support the development of more reliable digitization techniques.
By making PTB-Image publicly available and providing baseline digitization methods, we aim to encourage further advancements in ECG digitization research, ultimately improving access to historical ECG data and enabling broader applications in telemedicine and automated cardiac diagnostics.